### v2

# QUILL: Quotation Generation Enhancement of Large Language Models

Abstract
While Large language models (LLMs) have become excellent writing assistants, they still struggle with quotation generation. This is because they either hallucinate when providing factual quotations or fail to provide quotes that exceed human expectations. To bridge the gap, we systematically study how to evaluate and improve LLMs’ performance in quotation generation tasks. We first establish a holistic and automatic evaluation system for quotation generation task, which consists of five criteria each with corresponding automatic metric. To improve the LLMs’ quotation generation abilities, we construct a bilingual knowledge base that is broad in scope and rich in dimensions, containing up to 32,022 quotes. Moreover, guided by our critiria, we further design a quotation-specific metric to rerank the retrieved quotations from the knowledge base. Extensive experiments show that our metrics strongly correlate with human preferences. Existing LLMs struggle to generate desired quotes, but our quotation knowledge base and reranking metric help narrow this gap. Our dataset and code are publicly available at https://github.com/GraceXiaoo/QUILL.
QUILL: Quotation Generation Enhancement of Large Language Models
Jin Xiao1, Bowei Zhang1, Qianyu He2, Jiaqing Liang1thanks:  Corresponding author. Feng Wei3, Jinglei Chen3, Zujie Liang3, Deqing Yang1, Yanghua Xiao2 1School of Data Science, Fudan University 2Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University 3MYbank, Ant Group {jinxiao23, bwzhang24, qyhe21}@m.fudan.edu.cn, {liangjiaqing, yangdeqing, shawyh}@fudan.edu.cn {huodeng.wf, chenjinglei.cjl}@mybank.cn {jokieleung}@outlook.com
1 Introduction
Figure 1: An example of prevalent issues in Quotation Generation (QR) by LLMs. In QR tasks, LLMs often fabricate sentences, leading to quotation hallucination. Additionally, the generated quotes may not align with the context, resulting in contextual inconsistency and semantic incoherence. Finally, the sentences produced by LLMs tend to be overly common, resulting in a lack of novelty in quotations.
Figure 2: The framework for our Quotation Generation (QG) task research. We first establish an evaluation system with 5 evaluation criteria and automatic metrics, then build a quotation knowledge base covering multiple languages, topics and eras, and finally propose a quotation-specific reranking metric to rerank the quotations recalled in the RAG stage and improve the performance of QG tasks.
Famous quotations Tan et al. (2015a) are vital in academic and everyday communication. They lend authority to arguments and enhance persuasiveness, as they often stem from historically influential figures whose ideas have endured. Additionally, these quotations elevate the literary and artistic quality of a text, making discussions more engaging. They also facilitate comprehension of complex concepts, enabling readers to grasp ideas efficiently through concise expressions Vaswani et al. (2023).
The task of Quotation Generation (QG) seeks to produce suitable quotations to deepen the context in large language models (LLMs)  Anil et al. (2023); Achiam et al. (2023); Touvron et al. (2023). However, LLMs encounter significant challenges in this domain, as illustrated in Figure 1. Primarily, the generated quotes frequently fail to correspond to genuine famous quotations and are often inaccurately attributed, a phenomenon termed "Quotation Halluciantion." Huang et al. (2023); Bang et al. (2023); Guerreiro et al. (2023) Additionally, these quotes don’t align with the contextual meaning, resulting in a lack of coherence within the paragraph. Furthermore, LLMs exhibit a tendency to reproduce well-known quotes, which diminishes novelty and restricts creative expression.
Although the issues of Quote Generation task are particularly problematic in LLMs, there is currently no effective solutions. Previous studies Qi et al. (2022a) were based on representative pre-trained language models such as BERT Devlin et al. (2019), and it remains under-explored on the problem of quotation hallucination with LLMs. And there is currently no systematic and comprehensive benchmark to evaluate the quotation generation ability of LLMs.
To tackle these challenges, we introduce QUILL for QUotation GeneratIon enhancement of Large Language Models, a framework integrating an automatic evaluation system and an innovative and effective solution to improve quotation generation performance of LLMs.The framework of QUILL is shown in Fig. 2. QUILL presents a comprehensive benchmark comprising 7 quotation domains and 16 real-world scenarios to evaluate large models’ quotation generation abilities systematically, which consists of 5 highly interpretable and rigorous criteria with automatic evaluation metrics (Fig. 1): (1) Quotation Authenticity: Confirm whether the quoted quotes are real quotes from famous people to prevent misquotations or fabrications. (2) Quotation Credibility: Verify whether the quotation satisfies the author or source mentioned in the context (if any) to ensure the credibility of the quoted content. (3) Semantic Matching: Evaluate whether the semantics of the quoted quote align with the context. (4) Semantic Fluency: Evaluate the extent to which the cited quotation affects the fluency of the paragraph. (5) Quotation Novelty: Evaluate the degree of uniqueness of the quote.
Additionally, based on the task’s essential characteristics, we introduce an innovative Quotation-Specific Reranking Metric Karpukhin et al. (2020); Lewis et al. (2021); Chern et al. (2023) to improve model performance in QG tasks. To facilitate the task, we also established a comprehensive and high-quality knowledge database containing up to 32,022 quotes. This database spans both Chinese and English languages, various authors, different eras, and diverse topics, which ensures the wide applicability and generalization of our method. To the best of our knowledge, our work is the first systematic investigation into the automatic evaluation and enhancement of quotation generation performance in LLMs. To summarize, our contributions are mainly four-fold:
  1. 1.
    We establish a holistic and automatic evaluation system for the quotation generation task, consisting of five highly interpretable and rigorous criteria, facilitating both human and automatic evaluation of this task.
  2. 2.
    We construct a comprehensive and high-quality knowledge database containing up to 32,022 quotes, complete with authors or sources.
  3. 3.
    We design a fine-grained quotation-specific metric to rerank the retrieved quotations from the knowledge base.
  4. 4.
    We conduct extensive experiments to verify that our metrics are strongly correlate with human preference and significantly effective in both open-source and closed-source LLMs.
2 Related Work
2.1 Quotation
Previous research about quotation mainly focused on quote recommendation Tan et al. (2015a). The task of quote recommendation was initially proposed by Tan et al. (2015a). They proposed a learning ranking framework for the task, which integrates 16 manually crafted features. Lee et al. (2016) combined four different methods for recommending famous quotes, including matching granularity adjustment (a statistical context quote correlation prediction method), random forest, CNN, and LSTM. Wang et al. (2020) utilized an encoder-decoder framework to generate speech responses based on separate modeling of dialogue history and current query. Wang et al. (2021) used semantic matching to encode multi round dialogue histories using Transformer  Vaswani et al. (2023) and GRU  Cho et al. (2014), and encoded quotes using Transformer. However, previous studies do not take into account the quotation generation capabilities of the large models themselves, nor did they propose a systematic and comprehensive evaluation system or benchmark to assess model performance in scenarios involving famous quotes.
2.2 Hallucination
In the field of NLP, hallucinations typically refer to a phenomenon where generated content appears meaningless or does not align with the provided source Filippova (2020); Maynez et al. (2020). To address the issue of hallucinations in language models, two primary methods have been proposed: (1) preventing hallucinations during the training and generation processes, and (2) reducing hallucinations after generation. Manakul et al. (2023) introduced an alternative classification, dividing methods into black box and gray box approaches. Black box methods involve conducting factual checks without relying on external resources, either during or after generation. In contrast, gray box methods utilize external resources for validation. Other techniques for alleviating hallucinations include reranking generated sample responses  Dale et al. (2022) and improving beam search  Sridhar and Visser (2023). Recent mitigation technologies have also shown promise in reducing hallucinations  Mündler et al. (2024); Pfeiffer et al. (2023); Chen et al. (2023); Zhang et al. (2024); Agrawal et al. (2024). Although these methods have alleviated the quotation problem to a certain extent, they have not yet completely solved it, particularly in factual quotation and famous quotes.
3 Background
3.1 Task Formulation
Quotation Generation
Given a plain text $c=[t_{1},t_{2},\dots,t_{i},\dots,t_{n}]$, the goal of the Quotation Generation (QG) task is to generate quotes for the specified insertion point $i$. The left and right contexts, $c_{l}$ and $c_{r}$, are defined as $c_{l}=[t_{1},t_{2},\dots,t_{i}]$ and $c_{r}=[t_{i+1},\dots,t_{n}]$, respectively. In our work, we mainly focus on the ability of the model in quotation generation tasks.
Quotation Recommendation
In the Quotation Recommendation (QR) task, given the context $c=[t_{1},t_{2},\dots,t_{i},\dots,t_{n}]$, the objective is to select the most suitable quote from the given set $Q=\{q_{1},\dots,q_{|Q|}\}$ to insert at position $i$, where $q_{j}$ represents the $j$-th quote.
3.2 Preliminaries
Perplexity (PPL) is a crucial metric in natural language processing, reflecting a model’s predictive capability on text data and indicating the certainty of its next word prediction. Lower perplexity signifies greater confidence in the model’s predictions, demonstrating a stronger ability to generate or understand language. PPL of a language model given a sequence of words $w_{1},w_{2},\ldots,w_{N}$ is defined as:
$PPL\left(P_{r}\mid P_{l}\right)=\exp\left(-\frac{1}{s}\sum_{i=t+1}^{N}\log P(w_{i}\mid w_{1},\ldots,w_{i-1})\right)$ (1)
where $P_{l}$ is the given left paragraph, $P_{r}$ is the following context needs to be calculated, $P(w_{i}\mid w_{1},w_{2},\ldots,w_{i-1})$ is the probability of the word $w_{i}$ given its left context, and $s$ is equal to N-t+1, which is the length of the sequence in the following paragraph.
4 Evaluation System for QG
The accuracy and rationality of quoting famous quotes are crucial, as they directly affect the credibility and rigor of the content. Therefore, we establish a holistic and automatic evaluation system for QG task evaluation in LLMs, containing five criteria and further design automatic metrics for each criterion (Fig. 1).
Criteria
Considering the nature of the quotation task itself, we design the following five criteria: (1) Quotation Authenticity: Confirm whether the quoted quotes are real quotes from famous people to prevent misquotations or fabrications. (2) Quotation Credibility: Verify whether the quotation satisfies the author or source mentioned in the context (if any) to ensure the credibility of the quoted content. (3) Semantic Matching: Evaluate whether the semantics of the quoted quote align with the context. (4) Semantic Fluency: Evaluate whether the quoted quote affects the fluency of the original text. (5) Quotation Novelty: Evaluate the degree of uniqueness of the quote.
Evaluation Metrics
We propose automatic evaluation metrics for design standards, taking into account the essence of each metric. For any text containing the quote $q$, the segment preceding the quote is termed the left context $c_{l}$, while the segment following it is the right context $c_{r}$. The combination of these segments forms the speech context $c=[c_{l};c_{r}]$.
Quotation Authenticity. Authenticity of quotations is crucial as it ensures the reliability and credibility of information Kington et al. (2021). To verify the authenticity of the quoted celebrity quotes, we first search the quotation database for the information corresponding to the quote. If the database contains the information, we use the corresponding information to make a judgment. If not, we use different search engines (such as Google Scholar [^1] https://scholar.google.com/ and Baidu Scholar [^2] https://xueshu.baidu.com/) to recall the corresponding search results. Previous studies Han et al. (2024) have shown that GPT-4o OpenAI (2022) has excellent simple extraction capabilities, and the extraction task based on this study only has two fields, author and source. Therefore, we use GPT-4o to extract the corresponding field information, and then compare the results of different search engines. If the field information is different, manual comparison is required. For extraction details and validity of GPT-4o, please refer to Appendix C. Finally, based on the extracted information, we verify whether the quote genuinely originates from the specific celebrity. The final score is defined as follows:
$S_{a}=\begin{cases}1,&\text{if quote is real}\\ 0,&\text{if not real}\end{cases}$ (2)
Quotation Credibility. Generally speaking, in the context of quoting, the source of the quote will be mentioned, such as the author, classic literature, or other sources. Ensuring consistency between the citation and the mentioned author or source is crucial for maintaining contextual coherence and information accuracy. In order to confirm whether the citation meets the source restriction mentioned in the context, our study first extracts the source restriction of the context, and then compares and analyzes it with the extraction result of the previous indicator. If the source matches, the citation is marked as trustworthy, as shown in Fig.1. The final score for quotation credibility is defined as follows:
$S_{c}=\begin{cases}1,&\text{if restriction matches}\\ 0,&\text{if not match}\end{cases}$ (3)
Semantic Matching. Improper quotation may lead to misunderstandings or misinterpretations of the original meaning, thereby weakening the effectiveness and persuasiveness of the argument Quora (2020). Perplexity is a common metric in NLP, used to assess a language model’s predictive capability for text. Hence, we calculate the PPL of subsequent text based on a given prior text and quotation to evaluate the consistency between the quotation and its context. If the evaluation score is low, it implies that the citation aligns well with the following context in terms of semantics; otherwise, the rationality of the citation should be reconsidered. The formula is as follows:
$PPL_{m}=PPL\left(c_{r}\mid[c_{l};q]\right)$ (4)
where $c_{l}$ stands for the previous text, $q$ stands for the quotation, and $c_{r}$ stands for the following text.
To simplify computation, we normalize the PPL values to a range between 0 and 1. Given that a lower PPL indicates a higher degree of semantic alignment, we utilize an inverted Sigmoid function. The final calculation formula is as follows:
$S_{m}=\dfrac{1}{1+e^{k_{m}(PPL_{m}-\mu_{m})}}*100\%$ (5)
where $\mu_{m}$ represents the mean value of $PPL_{m}$, which is 35.243, and $k_{m}$ is determined using an empirical formula, yielding a value of 0.053. See the Appendix D for the specific calculation details.
Semantic Fluency. After quotation, it is necessary to ensure that the entire context is fluent and coherent to maintain semantic consistency and logical integrity Krumm et al. (2020). This study calculates the PPL of the entire context to measure the textual fluency of the overall context after inserting quotations. Lower perplexity indicates smoother overall contextual semantics. The calculation formula for semantic fluency is as follows:
$PPL_{f}=PPL_{q}\left([c_{l},q,c_{r}]\mid\cdot\right)$ (6)
where $c_{l}$ stands for the previous text, $q$ stands for the quotation, and $c_{r}$ stands for the following text.
Similarly, for normalizing the PPL values into a range from 0 to 1, the final score for semantic fluency is designed as follows:
$S_{f}=\dfrac{1}{1+e^{k_{f}(PPL_{f}-\mu_{f})}}*100\%$ (7)
where $\mu_{f}$ represents the mean value of $PPL_{f}$, which is 16.470, and $k_{f}$ is determined using an empirical formula, yielding a value of 0.500.
Quotation Novelty. Integrating novel quotations into established ideas enhances originality and personalizes the expression within academic discourse Savov (2021). To evaluate the extent to which the quote introduces new ideas or unique perspectives to the original context, we utilize the Bing [^3] https://www.bing.com/search engine to determine the number of Search Frequency corresponding to each quotation, applying a log10 transformation to quantify quotation popularity. In addition, to mitigate potential biases in search results, we also incorporate the quoted PPL value for supplementation. As a lower PPL indicates a higher frequency of text occurrence, it is inversely correlated with search frequency. Therefore, the formula is as follows:
$\text{Novelty}=\frac{PPL(q\mid\cdot)}{log_{10}(\text{Search Frequency})}$ (8)
where Search Frequency indicates the number of search results obtained by searching the quotation in the Bing search engine. In order to map the PPL value to a range of 0 to 1, since higher novelty means higher score, the positive sigmoid function is adopted here, and the final score is as follows:
$S_{n}=\dfrac{1}{1+e^{-k_{n}(Novelty-\mu_{n})}}*100\%$ (9)
where $\mu_{n}$ represents the mean value of Novelty, which is 10.67, and $k_{n}$ is determined using an empirical formula, yielding a value of 0.253.
5 Quotation Knowledge Base
5.1 Dataset Construction
In order to alleviate the problem of famous quote hallucination in LLMs, we develop a comprehensive bilingual and multi-topic quotation corpus designed to enhance retrieval quotation tasks during the RAG stage , as shown in Tab. 1. This corpus is structured into three distinct components: the English, the Standard Chinese, and the Classical Chinese. To improve the application scope and practical value of the corpus, we ensure comprehensive coverage of both common and specialized fields and also implement stringent quality control measures. Each quote is manually reviewed to ensure accuracy and relevance.
Category Samples AvgLen AvgSearchFreq AvgNovelty
English 16,393 16 2,823,499 6.8
Standard Chinese 7,519 42 150,011 6.3
Classical Chinese 8,110 14 19,017 5.0
Total 32,022 24 997,509 6.0
Table 1: The statistics of our knowledge base. For each category, the AvgLen, AvgSearchFreq and AvgNovelty denote the average of the length, the frequency of Bing Search engine, and the value of Quotation Novelty respectively.
English Corpus
To construct the English quotation corpus, we extract approximately 27,260 quotes covering different topics from the BrainyQuote [^4] https://www.brainyquote.com/, A-ZQuote [^5] https://www.azquotes.com/ and Goodreads [^6] https://www.goodreads.com/ websites, categorizing them by topic and author.
Classical Chinese Corpus
Considering the representativeness and novelty of the Chinese corpus, we first collect some famous citations from Gushiwen [^7] https://www.gushiwen.cn/. Subsequently, given the limited number of citations, we utilize LLM to conduct a meaningful selection of the collected poems from BaiduHanyu. For instance, the seven-character quatrains in Tang poetry can be divided into two citations. Furthermore, to enhance the generalization of themes, we employ LLM to summarize the topics of the quotes. Finally , we collect over 9,233 citations with its poems, author and topics, including various genres such as Tang poetry and Song lyrics.
Standard Chinese Corpus
Regarding the Standard Chinese quotation corpus, we gather 13,453 quotes from the Guzimi [^8] https://www.juzimi.com.cn/mingyan/ and Mingyancidian [^9] http://mingyan.juzicidian.com websites, similarly categorized by topic and author.
Dataset Evolution
For those collected from diverse websites, the corpus have two limitations: (1) Semantic redundancy: the semantics of different quotations are too similar, especially when a long quotation includes a shorter one. (2) Lengthy quotations: some quotations are excessively long. Hence, We first utilized the Jaccard Similarity coefficient to address the issue of semantic redundancy. Then we set a restriction on the length of the citations and remove the extreme values based on the quotation ppl metric. Additionally, to facilitate the subsequent rerank stage of retrieval-augmented generation (RAG), we also pre-calculate the novelty of the quotations in the database. The specific calculation is detailed in Equation (8). Finally, we obtain a higher-quality corpus exceeding 32,022 entries. The statistics of our knowledge dataset is as show in the Tab. 1.
5.2 Dataset Statistics
In this part, we compare the statistics of our dataset with existing quotation-related resources, as shown in Tab.2. In contrast, our dataset is the first to consider quotation novelty, encompassing a wide range of topics and numerous authors, while recording and annotating their sources. Additionally, we have expanded the scale of the quotation dataset, thereby broadening its application scenarios and significance.
5.3 Quality Assessment by Human
After constructing the dataset, we manually check its quality. For each component, we randomly select 100 quotes and engage three annotators to verify their validity. The annotators use search engines  [^10] https://www.bing.com/ to locate references and evaluate both the authenticity of the quotes and the accuracy of their attributed authors and sources. Only quotes that satisfy both criteria are included in the final dataset. The final results are determined through a majority voting process. In the English, Standard Chinese and Classical Chinese components, 99, 97 and 98 quotations respectively met the established criteria. These results confirm the high quality of the dataset, which is derived from trustworthy sources such as published books and reputable citation websites.
Resource Size Topic Author Multilingual Novelty
LRQW  Tan et al. (2015b) 3,158 822 762 N N
QRDW  Ahn et al. (2016) 1,200 - - N N
QuoteR  Qi et al. (2022b) 13,550 - - Y N
Ours 32,022 2,301 9,708 Y Y
Table 2: The statistics of our dataset with existing quotation-related resources. Multilingual refers to the inclusion of two or more languages, Y denotes Yes, and N denotes No.
6 Quotation-specific Reranking Metric
In our study we introduce a fine-grained and end-to-end RAG solution to improving model performance in quotation tasks through introducing a straightforward and interpretable quotation-specific rerank metric to select the optimal quotation.
When the user inputs the context to be inserted, we use semantic similarity to recall the top k most relevant quotes from the knowledge database. However, while similarity assesses the semantic relevance between the quotation and the context, the QG task necessitates a more comprehensive approach. It requires not only that the semantics of the quote align with the context but also that the paragraph maintains fluency and incorporates novel citations. To enhance the performance of LLMs in QG, we propose three evaluative sub-indicators as shown in Fig. 2:
Quotation Matching
Quotation matching emphasizes the completion of the quotation itself and its alignment with the subsequent text MacLaughlin and Smith (2021). This is accomplished by calculating the PPL of the remaining portion of the quotation, given the preceding text and the initial k characters of the quotation. Generally, lower PPL values suggest that the model produces more accurate and coherent quotations. The specific calculation formula is as follows:
$PPL_{q}=PPL\left([q_{n-t};c_{r}]\mid[c_{l};q_{t}]\right)$ (10)
where $n$ represents the length of the quote, $q_{t}$ represents the first $t$ characters of the quote, $q_{n-t}$ represents the remaining $n-t$ characters of the quote.
Semantic Matching
Semantic matching is concerned with ensuring semantic consistency and logical coherence within the context. This is achieved by predicting the PPL of the subsequent text, given the preceding text and the entire quote. A lower PPL value means that the quotation is more semantically consistent with the following context. The calculation formula is as Equation (4).
Figure 3: 7 common categories and 21 scenarios details of the evaluation dataset.
Novelty
The Novelty metric evaluates the originality of generated quotations. By avoiding repetition and clichés, this metric ensures that content remains fresh and engaging, providing unique perspectives across various contexts. The specific calculation formula is as Equation (8).
To integrate the advantages of the three indicators, we employ a weighted average method, utilizing it as our final quotation-specific rerank metric. This comprehensive indicator seeks to balance semantic matching, fluency, and novelty, thereby enhancing the overall quality of model-generated citations. Finally, after the rerank stage, we select the top-1 quote including author or source information, and add it to the prompt. Then, the model inserts and rewrites quotes dynamically in the context, and ultimately outputs the results we need.
7 Experiments
Model $S_{a}$ $S_{c}$ $S_{m}$ $S_{f}$ $S_{n}$ $Avg$
Chinese-oriented Models
ChatGLM3-6B 0.56 0.20 0.72 0.73 0.71 0.58
Qwen1.5-7B-Chat 0.63 0.15 0.72 0.68 0.71 0.58
Qwen1.5-14B-Chat 0.66 0.16 0.72 0.69 0.73 0.60
Qwen1.5-72B-Chat 0.72 0.16 0.71 0.71 0.67 0.60
Deepseek-R1 0.70 0.39 0.72 0.76 0.49 0.61
English-oriented Models
Mixture-7B-v0.2 0.77 0.08 0.70 0.74 0.55 0.57
Llama2-7B-Chat-hf 0.46 0.15 0.73 0.73 0.74 0.56
Llama2-13B-Chat-hf 0.48 0.15 0.74 0.72 0.74 0.56
Llama2-70B-Chat-hf 0.60 0.11 0.69 0.67 0.62 0.55
Close-source Models
GPT-3.5-turbo 0.62 0.11 0.71 0.72 0.62 0.56
GPT-4o 0.79 0.23 0.71 0.74 0.58 0.61
Ours 1.00 1.00 0.75 0.75 0.81 0.86
Table 3: Comparison of performance of various models on our evaluation system for QG tasks.
In this section, we conduct experiments to verify the effectiveness of our method and metrics.
7.1 Experiment Setup
Evaluation Dataset
In constructing the evaluation dataset, our study select 7 common categories: economy, diplomacy, journalism, academia, law, technology, and life. Additionally, 21 frequently cited scenarios are chosen to encompass various aspects of the knowledge system, as show in Fig. 3. To enhance the dataset’s diversity, standard Chinese, classical Chinese, and English texts are also included. Initially, we gather quotes from each scenario to ensure diversity, richness, and relevance to the selected fields. After collecting these quotes, they are used as keywords to search on major search engines like Google [^11] https://www.google.com/, Bing [^12] http://www.bing.com, and Baidu [^13] https://www.baidu.com/. Then articles containing these quotes are identified, and the relevant context is extracted. To guarantee the dataset’s quality, we perform preprocessing and cleaning, which involved removing duplicates, correcting errors, and eliminating ambiguities. Then, we conduct manual sampling and validation to evaluate and ensure the dataset’s quality and usability. Finally, we obtain the evaluation dataset that comprises 600 context-quote pairs.
Models
We evaluate 9 models ranging from their model sizes and structures, which fall into three categories: Chinese-oriented models, English-oriented models, and Close-source models.
Models for PPL Calculation
We employ two advanced language models, Qwen2-7B Bai et al. (2023) and Llama3-8B Touvron et al. (2023). These models are used to compute the PPL of the context given the preceding text. Subsequently, the average PPL values calculated by the two models are taken as final PPL values, which balances the judgments of the two models and reduces potential bias introduced by any single model. Since larger models tend to produce lower PPL for the same text, we recommend using the same PPL calculation models in this study when evaluating QG tasks.
Method HR@1 HR@3 nDCG@1 nDCG@3 MRR
Vanilla 0.13 0.43 0.50 0.72 0.35
Supervised
BM25 0.19 0.50 0.54 0.71 0.39
monoT5 (3B) 0.31 0.61 0.65 0.77 0.48
Unsupervised
UPR (FLAN-T5-XL) 0.31 0.52 0.63 0.74 0.46
bge-reranker-large 0.32 0.55 0.71 0.82 0.47
LLM API (Permutation Generation)
GPT-3.5-turbo 0.33 0.61 0.72 0.84 0.50
GPT-4o 0.43 0.63 0.74 0.88 0.55
Quotation-specific Reranking Metric
[…]
Table 4: Performance of different rerank metrics on Hit@1, Hit@3, nDCG@1, nDCG@3 and MRR. $PPL_{q}$, $PPL_{m}$ and Novelty are as defined in Section 6, and $PPL_{avg}$ is the average of $PPL_{q}$ and $PPL_{m}$. Best performing reranker method are marked bold.
Figure 4: Correlation between our automatic evaluation metrics and human ratings. To avoid overlapping points, random jitters sampled from $N(0,{0.05}^{2})$ were added to human ratings after fitting the regression.
Naive-0-Shot Naive-1-Shot Naive-2-Shot Naive-CoT
Model $S_{a}$ $S_{c}$ $S_{m}$ $S_{f}$ $S_{n}$ $Avg$ $S_{a}$ $S_{c}$ $S_{m}$ $S_{f}$ $S_{n}$ $Avg$ $S_{a}$ $S_{c}$ $S_{m}$ $S_{f}$ $S_{n}$ $Avg$ $S_{a}$ $S_{c}$ $S_{m}$ $S_{f}$ $S_{n}$ $Avg$
Chinese-oriented Models
ChatGLM3-6B 0.56 0.20 0.72 0.73 0.71 0.58 0.59 0.13 0.72 0.68 0.67 0.56 0.62 0.13 0.72 0.68 0.68 0.57 0.64 0.16 0.71 0.69 0.67 0.57
Qwen1.5-7B-Chat 0.63 0.15 0.72 0.68 0.71 0.58 0.66 0.13 0.72 0.70 0.71 0.59 0.67 0.13 0.71 0.70 0.69 0.58 0.67 0.13 0.72 0.69 0.69 0.59
Qwen1.5-14B-Chat 0.66 0.16 0.72 0.69 0.73 0.60 0.68 0.17 0.72 0.67 0.71 0.60 0.74 0.18 0.71 0.71 0.65 0.60 0.69 0.18 0.72 0.73 0.68 0.60
Qwen1.5-72B-Chat 0.72 0.16 0.71 0.71 0.67 0.60 0.67 0.21 0.72 0.72 0.67 0.60 0.63 0.18 0.72 0.71 0.65 0.58 0.78 0.20 0.70 0.71 0.65 0.61
Deepseek-R1 0.70 0.39 0.72 0.76 0.49 0.61 0.67 0.38 0.71 0.71 0.54 0.60 0.71 0.38 0.71 0.72 0.54 0.62 0.77 0.35 0.71 0.74 0.54 0.62
English-oriented Models
Mixture-7B-v0.2 0.77 0.08 0.70 0.74 0.55 0.57 0.82 0.17 0.71 0.75 0.52 0.59 0.82 0.15 0.70 0.75 0.46 0.58 0.77 0.09 0.71 0.73 0.58 0.58
Llama2-7B-Chat-hf 0.46 0.15 0.73 0.73 0.74 0.56 0.46 0.09 0.73 0.71 0.66 0.53 0.44 0.12 0.73 0.74 0.67 0.54 0.49 0.14 0.74 0.73 0.70 0.56
Llama2-13B-Chat-hf 0.48 0.15 0.74 0.72 0.74 0.56 0.44 0.10 0.74 0.72 0.74 0.56 0.50 0.13 0.73 0.68 0.74 0.57 0.45 0.10 0.73 0.67 0.74 0.55
Llama2-70B-Chat-hf 0.60 0.11 0.69 0.67 0.62 0.55 0.65 0.20 0.71 0.66 0.67 0.58 0.70 0.20 0.71 0.69 0.63 0.59 0.75 0.13 0.71 0.68 0.66 0.59
Close-source Models
GPT-3.5-turbo 0.62 0.11 0.71 0.72 0.62 0.56 0.72 0.16 0.71 0.75 0.59 0.59 0.73 0.14 0.71 0.74 0.57 0.58 0.76 0.10 0.71 0.70 0.58 0.57
GPT-4o 0.79 0.23 0.71 0.74 0.58 0.61 0.75 0.24 0.70 0.74 0.60 0.61 0.80 0.23 0.71 0.76 0.57 0.62 0.83 0.22 0.71 0.73 0.60 0.62
Table 5: Comparison of performance of various models on our evaluation system for QG tasks in in Naive-0-shot, Naive-1-shot, Naive-2-shot and Naive-cot settings. In these naive experimental setup, our experiment does not employ RAG or rerank metrics. Instead, it relies solely on a specifically designed prompt to enable the models to execute the QG task. The prompt for each setting is detailed in the Appendix  LABEL:naive.
7.2 Results
We conduct experiments on models of different ranges and sizes on our benchmark, and the results are shown in Tab. 3. For more detailed analysis, please refer to Appendix  E.
Severity of Quotaiton Hallucination
The results show that more than half of the citations generated by LLaMA2-13B-Chat are not genuine quotes. Furthermore, despite varying parameter sizes, all models demonstrate suboptimal performance on the QR task, especially on the $S_{c}$ metric. Even the best-performing model, GPT-4o, only scores 0.23 on the $S_{c}$ indicator, highlighting the critical need to address the quotation hallucination problem.
Performance of Quotation-specific Reranking Metric
Notably, our Quotation-specific Reranking method achieves the best results in each indicator, demonstrating the effectiveness of our method. Since our method retrieves the most relevant and appropriate citations from the quotation database, it ensures the authenticity and credibility of the citations. Therefore, both $S_{a}$ and $S_{c}$ are equal to 1. In addition, our method can effectively improve the novelty of citations and alleviate the problem of generating common citations with LLMs.
Comparison between Model Sizes
We conduct further analysis on different model sizes. Within the same series, larger models tend to show improved performance. This indicates that larger models have richer quotation memory and stronger instruction-following capabilities.
7.3 Ablation Study
Correlations with Human Ratings
We randomly select five samples for each scenario from the evaluation dataset, totaling 105 data samples. Due to the varying requirements for background knowledge across different categories (Fig.3), this study specifically invites expert professors in relevant fields to manually score evaluation metrics. Since Quotation Authenticity and Credibility are objective factual metrics, the manual evaluation primarily focused on the remaining three metrics. The process is independently conducted by experts, who are free to consult relevant literature and materials during the evaluation to ensure the reliability and objectivity of the results. Subsequently, we employ correlation analysis to assess the degree of association between various metrics and the overall evaluation results. As shown in Fig. 4, all metrics exhibit high levels of correlation. Specifically, the correlation coefficients are significantly higher than the threshold for statistical significance, indicating that our metric system effectively reflects the actual conditions of the evaluation subjects. For correlation analyses of specific categories, please refer to the Appendix F, where the results also reveal a significant correlation between manual and automated metrics for each category.
Correlation between Evaluation Metrics
We present the correlations among the five automatic metrics in Tab. 6. As shown, the correlations between the metrics are all weak. This indicates that the five metrics are mutually independent, making it necessary to evaluate each of them individually in order to obtain a comprehensive view of the citation generation task assessment.
Metric $S_{a}$ $S_{c}$ $S_{m}$ $S_{f}$ $S_{n}$
$S_{a}$ 1.000 -0.038 -0.018 -0.132 0.077
$S_{c}$0.038 1.000 -0.033 0.025 0.005
$S_{m}$0.018 -0.033 1.000 0.070 0.004
$S_{f}$0.132 0.025 0.070 1.000 0.002
$S_{n}$ 0.077 0.005 0.004 0.002 1.000
Table 6: Correlation Matrix between Evaluation Metrics
Effectiveness of Reranking Metrics
This study delves into the effectiveness of the rerank metric designed in our method and validates it through a series of ablation experiments. We adopt the following metrics: Hit Ratio at rank K (HR@K(K=1,3)), Normalized Discounted Cumulative Gain at rank K (NDCG@K(K=1,3)), and Mean Reciprocal Rank (MRR) for comparison. On our benchmark, we compare a range of defined quotation-rerank metrics with state-of-the-art supervised, unsupervised, and closed-source API-based reranking methods. The supervised baselines include: BM25 Nogueira and Cho (2019) and monoT5 Nogueira et al. (2020). The unsupervised baselines comprise UPR Sachan et al. (2023) and bge-reranker-large BAAI (2023). The closed-source API-based baselines include ChatGPT3.5 and ChatGPT4. As shown in Table 4, our simple yet effective quotation reranking metrics that demonstrate strong performance across various evaluation criteria. Notably, the $PPL_{\text{avg}}\text{+}\text{Novelty}$ metric excels among the four metrics and ranks just behind GPT-4 in the nDCG@3 metric. Importantly, both supervised and unsupervised methods underperform compared to our proposed reranking metrics. This indicates that our approach effectively captures the nuances of the quotation generation task, leading to superior citation recommendations.
Comparison between Prompt Strategies
We compare various prompting methods for QG tasks, including 0-shot, 1-shot, 2-shot, and Chain of Thought (CoT) Wei et al. (2023) strategies. For the CoT method, we implement a basic "let’s think step-by-step" approach. As shown in Table 5, among the four naive settings, the CoT method outperforms the others . The performance variations among the few-shot settings are not statistically significant, which suggests that the model’s in-context learning Dong et al. (2024) capability will not substantially enhance its quotation performance. In contrast, the logical reasoning stimulated by the CoT method improves the model’s quotation abilities to a certain degree.
Method Literal Sentence Recalled List Metric Rerank Human Rerank
BM25 Education empowers individuals to transform their lives and contribute to societal progress. [Q]. It fosters critical thinking, innovation, and social responsibility. By providing access to knowledge, education breaks down barriers and creates opportunities. It is a key driver of positive change and development.
Education is a human right with immense power to transform. On its foundation rest the cornerstones of freedom, democracy and sustainable human development. Education is the transmission of civilization. Education is the most powerful weapon which you can use to change the world
Education is the transmission of civilization Education is a human right with immense power to transform. On its foundation rest the cornerstones of freedom, democracy and sustainable human development Knowledge is power. Information is liberating. Education is the premise of progress, in every society, in every family.
Knowledge is power. Information is liberating. Education is the premise of progress, in every society, in every family Education is the most powerful weapon which you can use to change the world Education is a human right with immense power to transform. On its foundation rest the cornerstones of freedom, democracy and sustainable human development
Education is the most powerful weapon which you can use to change the world Knowledge is power. Information is liberating. Education is the premise of progress, in every society, in every family Education is the transmission of civilization
The function of education is to teach one to think intensively and to think critically. Intelligence plus character - that is the goal of true education The function of education is to teach one to think intensively and to think critically. Intelligence plus character - that is the goal of true education The function of education is to teach one to think intensively and to think critically. Intelligence plus character - that is the goal of true education
Ours Education empowers individuals to transform their lives and contribute to societal progress. [Q]. It fosters critical thinking, innovation, and social responsibility. By providing access to knowledge, education breaks down barriers and creates opportunities. It is a key driver of positive change and development.
Education is a human right with immense power to transform. On its foundation rest the cornerstones of freedom, democracy and sustainable human development Education is the most powerful weapon which you can use to change the world Education is the most powerful weapon which you can use to change the world
Education is the transmission of civilization Education is a human right with immense power to transform. On its foundation rest the cornerstones of freedom, democracy and sustainable human development Knowledge is power. Information is liberating. Education is the premise of progress, in every society, in every family
Knowledge is power. Information is liberating. Education is the premise of progress, in every society, in every family Knowledge is power. Information is liberating. Education is the premise of progress, in every society, in every family Education is a human right with immense power to transform. On its foundation rest the cornerstones of freedom, democracy and sustainable human development
Education is the most powerful weapon which you can use to change the world Education is the transmission of civilization Education is the transmission of civilization
The function of education is to teach one to think intensively and to think critically. Intelligence plus character - that is the goal of true education The function of education is to teach one to think intensively and to think critically. Intelligence plus character - that is the goal of true education The function of education is to teach one to think intensively and to think critically. Intelligence plus character - that is the goal of true education
Table 7: The examples of recalled candidates reranked via different rerank metrics and human evaluation. The indicators [Q] denotes the insertion positions of the given context. A darker shade of green indicates a higher rank bestowed by humans. See the Appendix for a detailed comparison of the unsupervised UPR, the closed- source model GPT-3.5-turbo, and our approach.
7.4 QUILL Application
In this study, we conduct a comprehensive case analysis to demonstrate the efficacy and alignment of our reranking metric with human evaluations. As illustrated in Tab. 7, we focus on several key models for comparison: the supervised BM25 and our own reranking metric, which combines average perplexity ($\text{PPL}_{avg}$) with novelty. Additionally, we manually sort and annotate the top-5 quote list initially recalled, serving as a benchmark for comparison. The findings reveal that our metric exhibits a higher correlation with human sorting than the other methods, underscoring its broad applicability and effectiveness. See the Appendix for a detailed comparison of the unsupervised UPR, the closed-source model GPT-3.5-turbo, and our approach.
8 Conclusion
In this paper, we systematically explore methods to enhance the performance of quotation generation tasks in LLMs. Initially, we establish a holistic and automatic evaluation system consisting of five highly interpretable and rigorous criteria , facilitating both human and automatic evaluation of this task. Then, we construct a comprehensive and high-quality knowledge database containing up to 32,022 quotes, complete with authors or sources. Moreover, we design a fine-grained quotation-specific metric to rerank the retrieved quotations from the knowledge base to improve QG performance. Additionally, we conduct extensive experiments to verify that our metrics are strongly correlate with human preference and significantly effective in both open-source and closed-source LLMs.
Limitations
This study highlights several limitations. We primarily use Perplexity (PPL) to evaluate text fluency. Although PPL is widely applied, it only measures the divergence between the model’s and true probability distributions. Future research should integrate additional metrics or human evaluations for a more comprehensive assessment. Additionally, our analysis is restricted to specific contexts with clear correlations before and after quoted content. While informative, this approach does not cover a wide range of quoting scenarios. Future studies should explore diverse applications for more generalizable insights.
Appendix
Appendix A Details of Evaluation Dataset
We also conducted manual analysis on the Evaluation Dataset, selecting 275 quotes from numerous context-quote pairs, dividing into Chinese and English, which categories and scenarios details are shown in Figure.3. After statistics, there are 204 Chinese samples and 71 English samples, with a total of 144 Chinese and English authors.
Appendix B Details of Quotation Knowledge Base
This chapter further analyzes the data details in the quotation corpus, which is divided into three languages: English, Standard Chinese, and Classical Chinese, all classified by topic and author. The number of topics and authors for each language is shown in Table.8. In addition, we also conduct analysis on the proportion of different topics in each language in the corpus, as shown in Figure. 56. for specific topics and proportions.
Language Type Topic Author Total
English 1,216 6,624 16,393
Standard Chinese 228 2,377 7,519
Classic Chinese 869 876 8,110
Table 8: The specific topics, authors, and total count of the quotation corpus.
Appendix C Effectiveness of GPT-4o Extraction
Previous studies have demonstrated that GPT-4o exhibits superior performance in simple information extraction tasks under zero-shot settings Han et al. (2024). In our context, we primarily extract authors and sources, which involves only two fields, thus categorizing this as a simple information extraction task. Therefore, we believe that GPT-4o is capable of achieving excellent extraction performance. We also conduct experiments to validate the extraction effectiveness of GPT-4o in our task. We extract 100 citations containing authors or sources from a citation knowledge base in three different languages and use these as keywords to search on the Bing search engine. In the returned search results, GPT-4o is utilized to extract the authors or sources of the citations. Ultimately, we assess the matching degree between the fields extracted by GPT-4o and the annotated fields in the knowledge base. The specific results are as follows:
Language Type Extraction accuracy
English 97%
Standard Chinese 95%
Classic Chinese 98%
Table 9: Extraction and verification results of ChatGPT
Appendix D Details of the Inverted Sigmoid Function
To map the calculated Perplexity (PPL) values to a range of [0, 1], this study employs the Sigmoid function, which not only maps the scores to [0, 1] but also handles positive extreme values in the data. For the two key parameters of the Sigmoid function, $k$ and $\mu$, the calculation methods used in this study are as follows:
For $\mu$: The Sigmoid function outputs 0.5 when $x=\mu$, where the slope is at its maximum. Typically, $\mu$ is set to the median or mean of the data, ensuring that the middle values are mapped to 0.5. In this study, we choose the mean of the data as the value for $\mu$.
For $k$: The slope parameter $k$ controls the "compression degree" of the mapping. A larger $k$ results in a steeper Sigmoid curve, which is suitable for data with a concentrated distribution. In contrast, a smaller $k$ results in a gentler curve, making it more appropriate for data with a wide range or extreme outliers. This study calculates $k$ based on an empirical formula as follows:
$k=\dfrac{ln(9)}{Q_{95}-Q_{5}}$ (11)
where $ln(9)\approx$ 2.2, corresponding to the span of the Sigmoid function from 0.1 to 0.9, $Q_{5}$ represents the 5% digit of the data, and $Q_{95}$ represents the 95% digit of the data.
Appendix E More Analysis of Experimental Results
Due to space limitations, we provide more experimental results analysis in this section.
Comparison between Model Types
The performance comparison between the Chinese-oriented group and the English-oriented group on the Chinese-English benchmark reveals no significant differences, suggesting that the model’s quotation ability is not language-dependent. Overall, the current opensource small to large-scale models exhibit a relatively small performance gap compared to close-source models, indicating the universality of the issue of quotation hallucination in LLMs.
Appendix F Details of Human Evaluation Metrics
We randomly selected 5 samples for each scenario from the evaluation dataset, totaling 105 data. Since different scenarios have different requirements for background knowledge, this study specially invited professional professors in related fields to manually score different categories of data. The scoring process was completed independently by experts, and relevant literature and materials were freely available during the review process to ensure the reliability and objectivity of the scoring results. In addition, we also further analyzed the correlation analysis of each specific category. The results are shown in the Table. 10. It can also be seen from the results that the manual indicators and automatic indicators of each category are also significantly correlated.
Metric Overall Science Business News Academic Life Law Diplomacy
Authenticity 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Credibility 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Matching 0.75 0.74 0.72 0.74 0.67 0.70 0.83 0.71
Fluency 0.72 0.71 0.69 0.71 0.64 0.67 0.80 0.68
Novelty 0.71 0.70 0.68 0.70 0.63 0.66 0.79 0.67
Table 10: Metric evaluation results across different categories.
Appendix G Details of Naive Setting Prompts
For the naive experimental settings, we also disclose its prompt in detail, see Table.12 for Naive-0-Shot, Table.13 for Naive-1-Shot, and Table.14 for Naive-Cot setting.
Appendix H More Cases of QUILL Application
In this study, we conduct a comprehensive case analysis to demonstrate the efficacy and alignment of our reranking metric of the unsupervised UPR, the closed-source model GPT-3.5-turbo, and our approach with human evaluation. As shown in Table.11, we show the results of rag without rearrangement, index rearrangement and manual evaluation. The darker the color, the higher the manual evaluation score.
Figure 5: The specific topic distribution of the English quotation corpus.
Figure 6: The specific topic distribution of the Classic Chinese quotation corpus.
Method Literal Sentence Recalled List Metric Rerank Human Rerank
GPT {CJK}UTF8gbsn 康托尔在研究集合论的过程中,深刻认识到数学探索的本质。他认为:[Q]。这句话强调了在数学研究中,发现和提出新的问题比解决已有问题更为重要。康托尔的集合论突破了传统数学的界限,提出了无穷集合和基数的概念,为数学理论的发展开辟了新的道路。
{CJK}UTF8gbsn 数学是科学的皇后,而数论是数学的皇后 {CJK}UTF8gbsn 新的数学方法和概念,常常比解决数学问题本身更重要 {CJK}UTF8gbsn 在数学的领域中,提出问题的艺术比解答问题的艺术更为重要
{CJK}UTF8gbsn 新的数学方法和概念,常常比解决数学问题本身更重要 {CJK}UTF8gbsn 在数学的领域中,提出问题的艺术比解答问题的艺术更为重要 {CJK}UTF8gbsn 新的数学方法和概念,常常比解决数学问题本身更重要
{CJK}UTF8gbsn 在数学的领域中,提出问题的艺术比解答问题的艺术更为重要 {CJK}UTF8gbsn 数学是一种理性的精神,使人类的思维得以运用到最完善的程度 {CJK}UTF8gbsn 数学是一种理性的精神,使人类的思维得以运用到最完善的程度
{CJK}UTF8gbsn 数学是一种理性的精神,使人类的思维得以运用到最完善的程度 {CJK}UTF8gbsn 数学之所以有高声誉,另一个理由就是数学使得自然科学实现定理化,给予自然科学某种程度的可靠性 {CJK}UTF8gbsn 数学之所以有高声誉,另一个理由就是数学使得自然科学实现定理化,给予自然科学某种程度的可靠性
{CJK}UTF8gbsn 数学之所以有高声誉,另一个理由就是数学使得自然科学实现定理化,给予自然科学某种程度的可靠性 {CJK}UTF8gbsn 数学是科学的皇后,而数论是数学的皇后 {CJK}UTF8gbsn 数学是科学的皇后,而数论是数学的皇后
Ours {CJK}UTF8gbsn 康托尔在研究集合论的过程中,深刻认识到数学探索的本质。他认为:[Q]。这句话强调了在数学研究中,发现和提出新的问题比解决已有问题更为重要。康托尔的集合论突破了传统数学的界限,提出了无穷集合和基数的概念,为数学理论的发展开辟了新的道路。
{CJK}UTF8gbsn 数学是科学的皇后,而数论是数学的皇后 {CJK}UTF8gbsn 礼义以生利,政事以成义 {CJK}UTF8gbsn 在数学的领域中,提出问题的艺术比解答问题的艺术更为重要
{CJK}UTF8gbsn 新的数学方法和概念,常常比解决数学问题本身更重要 {CJK}UTF8gbsn 数学是科学的皇后,而数论是数学的皇后 {CJK}UTF8gbsn 新的数学方法和概念,常常比解决数学问题本身更重要
{CJK}UTF8gbsn 在数学的领域中,提出问题的艺术比解答问题的艺术更为重要 {CJK}UTF8gbsn 数学之所以有高声誉,另一个理由就是数学使得自然科学实现定理化,给予自然科学某种程度的可靠性 {CJK}UTF8gbsn 数学是一种理性的精神,使人类的思维得以运用到最完善的程度
{CJK}UTF8gbsn 数学是一种理性的精神,使人类的思维得以运用到最完善的程度 {CJK}UTF8gbsn 数学是一种理性的精神,使人类的思维得以运用到最完善的程度 {CJK}UTF8gbsn 数学之所以有高声誉,另一个理由就是数学使得自然科学实现定理化,给予自然科学某种程度的可靠性
{CJK}UTF8gbsn 数学之所以有高声誉,另一个理由就是数学使得自然科学实现定理化,给予自然科学某种程度的可靠性 {CJK}UTF8gbsn 在数学的领域中,提出问题的艺术比解答问题的艺术更为重要 {CJK}UTF8gbsn 数学是科学的皇后,而数论是数学的皇后
UPR {CJK}UTF8gbsn 在荀子的政治哲学中,[Q]。是一个重要的命题。 礼义不仅是个人行为的规范,也是国家治理的基础。通过礼义,可以实现经济利益的最大化;通过公正的政务,可以实现社会正义。 荀子的这一思想在中国古代政治理论中占有重要地位,对后世的治国理政产生了深远影响。
{CJK}UTF8gbsn 故不积跬步,无以至千里;不积小流,无以成江海 {CJK}UTF8gbsn言无常信,行无常贞,惟利所在,无所不倾,若是则可谓小人矣 {CJK}UTF8gbsn礼义以生利,政事以成义
{CJK}UTF8gbsn 倘能生存,我当然仍要学习 {CJK}UTF8gbsn 礼义以生利,政事以成义 {CJK}UTF8gbsn 言无常信,行无常贞,惟利所在,无所不倾,若是则可谓小人矣
{CJK}UTF8gbsn 玉石不经雕琢,就不能成为有用的器物;人如果不学习,就不会明白道理 {CJK}UTF8gbsn 玉石不经雕琢,就不能成为有用的器物;人如果不学习,就不会明白道理 {CJK}UTF8gbsn 故不积跬步,无以至千里;不积小流,无以成江海
{CJK}UTF8gbsn 言无常信,行无常贞,惟利所在,无所不倾,若是则可谓小人矣 {CJK}UTF8gbsn 故不积跬步,无以至千里;不积小流,无以成江海 {CJK}UTF8gbsn 玉石不经雕琢,就不能成为有用的器物;人如果不学习,就不会明白道理
{CJK}UTF8gbsn 礼义以生利,政事以成义 {CJK}UTF8gbsn 倘能生存,我当然仍要学习 {CJK}UTF8gbsn 倘能生存,我当然仍要学习
Ours {CJK}UTF8gbsn 在荀子的政治哲学中,[Q]。是一个重要的命题。 礼义不仅是个人行为的规范,也是国家治理的基础。通过礼义,可以实现经济利益的最大化;通过公正的政务,可以实现社会正义。 荀子的这一思想在中国古代政治理论中占有重要地位,对后世的治国理政产生了深远影响。
{CJK}UTF8gbsn 故不积跬步,无以至千里;不积小流,无以成江海 {CJK}UTF8gbsn 礼义以生利,政事以成义 {CJK}UTF8gbsn 礼义以生利,政事以成义
{CJK}UTF8gbsn 倘能生存,我当然仍要学习 {CJK}UTF8gbsn 言无常信,行无常贞,惟利所在,无所不倾,若是则可谓小人矣 {CJK}UTF8gbsn 言无常信,行无常贞,惟利所在,无所不倾,若是则可谓小人矣
{CJK}UTF8gbsn 玉石不经雕琢,就不能成为有用的器物;人如果不学习,就不会明白道理 {CJK}UTF8gbsn 玉石不经雕琢,就不能成为有用的器物;人如果不学习,就不会明白道理 {CJK}UTF8gbsn 故不积跬步,无以至千里;不积小流,无以成江海
{CJK}UTF8gbsn 言无常信,行无常贞,惟利所在,无所不倾,若是则可谓小人矣 {CJK}UTF8gbsn 故不积跬步,无以至千里;不积小流,无以成江海 {CJK}UTF8gbsn 玉石不经雕琢,就不能成为有用的器物;人如果不学习,就不会明白道理
{CJK}UTF8gbsn 礼义以生利,政事以成义 {CJK}UTF8gbsn 倘能生存,我当然仍要学习 {CJK}UTF8gbsn 倘能生存,我当然仍要学习
Table 11: Additional example of recalled candidates reranked via different rerank metrics and human evaluation. The indicators [Q] denotes the insertion positions of the given context. A darker shade of green indicates a higher rank bestowed by humans.
/* Task prompt */
Suppose you are a literary scholar and are familiar with many famous people’s quotes. You are required to populate contextualised quotes based on user input text within the specified [Q] symbols.
/* Output requirements */
1. The famous quotes must be quotes from a famous person in history or in the present, Please output the quote in English.
2. The quote should be closely related to the context, so that the context is more reasonable, smooth and beautiful.
3. If there is a specified author in the context, the famous quote must be given according to the corresponding restrictions.
4. Output Formate: "quote".
5. Only output the quote, NO MORE INFORMATION!
6. The number of quote should be 5 to 30 words.
/* Input */
—INPUT—
{Query}
—OUTPUT—
Table 12: The details of the prompt for Naive-0-Shot setting.
/* Task prompt */
Suppose you are a literary scholar and are familiar with many famous people’s quotes. You are required to populate contextualised quotes based on user input text within the specified [Q] symbols.
/* Output requirements */
1. The famous quotes must be quotes from a famous person in history or in the present, Please output the quote in English.
2. The quote should be closely related to the context, so that the context is more reasonable, smooth and beautiful.
3. If there is a specified author in the context, the famous quote must be given according to the corresponding restrictions.
4. Output Formate: "quote".
5. Only output the quote, NO MORE INFORMATION!
6. The number of quote should be 5 to 30 words.
/* Example */
—INPUT—
.[Q], said by Confucius in Analects of Confucius - Wei Linggong. So is reading. Hard reading is the foundation, good reading is the key. In order to read effectively, you also need to make use of its "tools".
—OUTPUT—
"To do a good job, you must first sharpen your tools."
/* Input */
—INPUT—
{Query}
—OUTPUT—
Table 13: The details of the prompt for Naive-1-Shot setting.
/* Task prompt */
Suppose you are a literary scholar and are familiar with many famous people’s quotes. You are required to populate contextualised quotes based on user input text within the specified [Q] symbols.
/* Output requirements */
1. The famous quotes must be quotes from a famous person in history or in the present, Please output the quote in English.
2. The quote should be closely related to the context, so that the context is more reasonable, smooth and beautiful.
3. If there is a specified author in the context, the famous quote must be given according to the corresponding restrictions.
4. Output Formate: "quote".
5. Only output the quote, NO MORE INFORMATION!
6. The number of quote should be 5 to 30 words.
Please think step by step then return the result!!!
/* Examples */
1: —INPUT—
.[Q], said by Confucius in Analects of Confucius - Wei Linggong. So is reading. Hard reading is the foundation, good reading is the key. In order to read effectively, you also need to make use of its "tools".
—OUTPUT—
"To do a good job, you must first sharpen your tools."
2: —INPUT—
.[Q]. As an ancient civilisation and a responsible power, it has always been China’s pursuit to help the world. By guiding the direction of the world’s changing circumstances with Chinese concepts, Chinese-style modernisation will advance and expand in benign interaction with the world, and will also strengthen the power for world peace and provide opportunities for the development of all countries.
—OUTPUT—
"Already wanting to be established, we should be established; already wanting to achieve, we should achieve."
/* Input */
—INPUT—
{Query}
—OUTPUT—
Table 14: The details of the prompt for Naive-CoT setting.
Appendix I nDCG Formulation
In our experiment, in order to get the relevance between quote and query, we first use GPT-4o to score the relevance and get the complete relevance list after manual sampling. Hence, given $m$ candidate quotes $Q=\{q_{1},q_{2},\cdots,q_{m}\}$, the nDCG@k is defined as follows:
$\text{nDCG}(k)=\frac{\text{DCG}(O_{\text{real}},k)}{\text{DCG}(O_{\text{ideal}},k)}$ (12)
$\text{DCG}(O,k)=\sum_{i=1}^{k}\frac{Rel_{i}}{\log_{2}(1+i)}$ (13)
where $O_{\text{ideal}}$ and $O_{\text{real}}$ represent the score list given by the ideal ranking relevance and the real ranking relevance respectively, $Rel_{i}$ denote the relevance score of the quote $q_{i}$.